OpenStudy (anonymous):

What two numbers when added together give you -4/5?

OpenStudy (asnaseer):

just select any one number, say 10, and call the other one x, then write an equation to represent the conditions, i.e: 10 + x = -4/5 and solve to find x.

OpenStudy (anonymous):

But that only gives me one number -54/5

OpenStudy (apoorvk):

the other no. is 10 ofcourse!!!

OpenStudy (asnaseer):

the other one is the one we picked at random - i.e. 10

OpenStudy (apoorvk):

the one you assumed in the first place!

OpenStudy (anonymous):

Okay so my problem is -1/5x2-2x+4 and I factored this to -1/5x2+10x-54/5x+4. How do I factor out -1/5x2+10x and -54/5x+4?

OpenStudy (asnaseer):

I don't understand what you are trying to do?

OpenStudy (anonymous):

I need to factor -1/5x^2-2x+4 so I get an answer in ( - )^2 form.

OpenStudy (asnaseer):

the equation you have given cannot be factored into something squared. it can, however, be factored into something squared minus something else.

OpenStudy (asnaseer):

are you trying to solve a quadratic by completing the square?

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

ok, so the actual problem you are given is:\[-\frac{1}{5}x^2-2x+4=0\]if this is correct, then I would first simplify it by multiplying both sides by 5 to get rid of the fraction

OpenStudy (anonymous):

f(x)=-1/5x2-2x+1/4 so I multiply both sides by 20?

OpenStudy (asnaseer):

hang on - you have just changed the equation - is the last term "4" or "1/4"?

OpenStudy (anonymous):

Sorry it is 1/4

OpenStudy (asnaseer):

and what exactly does the question ask you to do?

OpenStudy (anonymous):

Write in standard form of y(x)=a(x+h)^2+k... my notes say to get this you use the complete the square method. Is this correct?

OpenStudy (asnaseer):

yes - but it is not what I originally thought was the question, which was to find the roots of the equation. so here what you should do is notice first that the coefficient of x in a(x+h)^2+k is one. therefore you need to get your equation in that form. you can do this as follows:\[f(x)=-\frac{1}{5}x^2-2x+\frac{1}{4}=-\frac{1}{5}(x^2+10x-\frac{5}{4})\]

OpenStudy (asnaseer):

then, since you need to get this into (x+h)^2, look at just the terms involving x^2 and x - here you have \(x^2+10x\), and see if you can write this in the form(x+h)^2 (ignoring any final constant). try this and let me know what you get.

OpenStudy (asnaseer):

i.e. try to find an h such that:\[(x+h)^2=x^2+10x+something\]

OpenStudy (anonymous):

I'm sorry... how did you get the 10x?

OpenStudy (asnaseer):

We are trying to get \(x^2+10x\) into a form \((x+h)^2\) because the equation for f(x) above involved these two as the first two terms, i.e.:\[f(x)=-\frac{1}{5}(x^2+10x-\frac{5}{4})\]

OpenStudy (anonymous):

10x came from the -2x?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

\[-\frac{1}{5}\times10=-2\]

OpenStudy (asnaseer):

I took \(-\frac{1}{5}\) out as common factor of the whole equation.

OpenStudy (anonymous):

Okay so I get -1/5 (x^2+10x)=1/4.

OpenStudy (asnaseer):

no

OpenStudy (anonymous):

Sorry not =1/4 +1/4

OpenStudy (asnaseer):

I think you are getting confused with this, let me try and start again...

OpenStudy (asnaseer):

your original equation was given as:\[f(x)=-\frac{1}{5}x^2-2x+\frac{1}{4}\] agreed?

OpenStudy (anonymous):

Correct.

OpenStudy (asnaseer):

ok, so next I took the -1/5 out as a common factor to get:\[f(x)=-\frac{1}{5}(x^2+10-\frac{5}{4})\]understand so far?

OpenStudy (asnaseer):

sorry I missed the x in 10x

OpenStudy (asnaseer):

\[f(x)=-\frac{1}{5}(x^2+10x-\frac{5}{4})\]

OpenStudy (asnaseer):

with me so far?

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

ok, so next we see that within the brackets we have \(x^2+10x-\frac{5}{4}\). and we know we need to get this in form \((x+h)^2\). make sense?

OpenStudy (anonymous):

This is where I get lost... what number is h?

OpenStudy (asnaseer):

h is not known yet - that is what we need to find

OpenStudy (anonymous):

And our x is 10 correct?

OpenStudy (asnaseer):

no, x is an unknown of the equation

OpenStudy (anonymous):

Okay so this is what x and h are what we are looking for next.

OpenStudy (asnaseer):

all we are trying to do is rewrite \(x^2+10x-\frac{5}{4}\) in the form \((x+h)^2\). so we need o find a suitable value for h. so what we do is first ignore the \(-\frac{5}{4}\)constant term, and just concentrate on \(x^2+10x\). so we want to find a value for h such that:\[(x+h)^2=x^2+10x+something\]and we don't care what the \(something\) is.

OpenStudy (asnaseer):

so, can you find a suitable value for h such that when we expand \((x+h)^2\) we get 10 as the coefficient of x?

OpenStudy (asnaseer):

\[(x+h)^2=x^2+2hx+h^2\]so we need 2h=10, which means h must be 5. understand?

OpenStudy (anonymous):

Yes because were diving 2h/2=10/2 gives us 5.

OpenStudy (anonymous):

dividing

OpenStudy (asnaseer):

ok, so now we know h, we can get back to the original equation within the brackets, which was:\[x^2+10x-\frac{5}{4}\]and we know:\[(x+5)^2=x^2+10x+25\]therefore, we can write:\[x^2+10x-\frac{5}{4}=(x+5)^2-25-\frac{5}{4}\]

OpenStudy (asnaseer):

the "-25" is there to get rid of the "+25" from the expansion of (x+5)^2

OpenStudy (asnaseer):

agreed?

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

ok, so finally we plug this into the original equation to get:\[\begin{align} f(x)&=-\frac{1}{5}(x^2+10x-\frac{5}{4})\\ &=-\frac{1}{5}((x+5)^2-25-\frac{5}{4})\\ &=-\frac{1}{5}(x+5)^2+5+\frac{1}{4}\\ &=-\frac{1}{5}(x+5)^2+5\frac{1}{4}\\ \end{align}\]

OpenStudy (asnaseer):

and this is in the desired form of f(x)=a(x+h)^2+k

OpenStudy (asnaseer):

I hope I explained it well enough for you to understand.

OpenStudy (anonymous):

Wow. I was totally going to wrong way before. Thank you so much... This helps a lot. My teacher didnt explain the standard form in fractions so I was trying to go about it in another direction. I should be able to follow this example on the rest of my problems. Thanks again!

OpenStudy (asnaseer):

yw - I'm glad I was able to help.

OpenStudy (anonymous):

Will my axis of symmetry be 5, 5 1/4?